# Numerical study of hypershadows in higher–dimensional black holes

### v2

@@ 3.2 Myers-Perry Black Hole
In 1986, Myers and Perry derived a stationary, axisymmetric, and asymptotically flat vacuum solution as a generalization of the Kerr black hole to higher dimensions[Myers:1986un]. The 5D Myers–Perry solution is given by the line element:
@@
The Black hole mass is controlled by the same parameter $\mu$ introduced in Eq 3.2, and its two angular momenta are
@@
while $a$ and $b$ are the spin parameter associated with two independent rotation planes. These parameters are bounded by the condition $\mu\geq|a|+|b|$, which ensures the existence of an event horizon. The angular coordinates $\phi$ and $\psi$ range over the interval $0,2\pi]$, while $\theta$ takes values in $0,\pi/2]$.
@@
The metric (3.5) is invariant under the following transformation
@@
$\cos\bar{\theta}\,\partial_{\bar{\theta}}-\cot\bar{\theta}\sin\bar{\phi}\,\partial_{\bar{\phi}}+\frac{\sin\bar{\phi}}{\sin\bar{\theta}}\,\partial_{\bar{\psi}},$ (3.10)
@@
$\sin\bar{\phi}\,\partial_{\bar{\theta}}-\cot\bar{\theta}\cos\bar{\phi}\,\partial_{\bar{\phi}}+\frac{\cos\bar{\phi}}{\sin\bar{\theta}}\,\partial_{\bar{\psi}}.$ (3.11)
@@
3.2.1 Cohomogeneity–one Solution: $a=b$
The shadow of the five-dimensional Myers–Perry black hole was first investigated through two-dimensional projection images by [Papnoi:2014aaa]. Recently, the work [Novo:2024wyn] provides an analytical description of the hypershadow boundary in the cohomogeneity–one case, with $a=b$. We compute the hypershadow numerically, investigate its physical features, and verify the consistency with the analytical results.
In order to study the cohomogeneity–one Myers–Perry spacetime ($a=b$), the line element can be written as [myers2011myers]:
@@ 3.2.1 Cohomogeneity–one Solution: $a=b$
By evolving the geodesic equations[Frolov:2003en, Diemer:2014lba] derived from the Hamiltonian formalism under specified initial conditions, we reconstruct the 3D hypershadow using the backward ray–tracing method.
@@
Figure 4: Hypershadows of the Myers–Perry black hole ($a=b=0.45$) rendered with different $N_{\rm pix}$ and viewpoints. And the observer position is located at $\{65,\pi/2-1/20\}$.
To probe the geometric structure of the hypershadow, we fix the observation angles $\{\alpha,\beta,\zeta\}$ to zero and compute cross–sectional slices through its center along each coordinate plane, in order to examine and discuss their symmetries.
@@
[…]
@@
In this paper [Novo:2024wyn], the authors proposed a simpler parameterization of the MP hypershadow based on analytical construction. They concluded that the shape of the hypershadow is independent of the observer inclination angle $\theta_{\rm o}$. Varying $\theta_{\rm o}$ only results in a rigid rotation of the hypershadow in the image space, while the overall structure remains unchanged. Moreover, the geometry exhibits continuous rotational symmetry about the central axis. When $\theta_{\rm o}=\pi/2$, this central axis aligns with the X-axis.
@@
[…]
From Fig. 6, we find that as $\theta_{\rm o}\to\pi/2$, the difference between the original slice and its vertical reflection diminishes significantly. This suggests that the $XY$ central slice becomes symmetric under vertical reflection when $\theta_{\rm o}=\pi/2$. Consequently, this implies the existence of continuous rotational symmetry about the central axis. After confirming the approximate vertical symmetry near $\theta_{\rm o}\approx\pi/2$, we proceed to analyze the central slices in the $XY$ plane for various observer inclination angles, in order to further illustrate the rotational behavior of the hypershadow.
@@
Figure 7: XYXY central slices with different $\theta_{\rm o}$
From Fig. 7, we could get that when we decrease the value of $\theta_{\rm o}$, the central axis will counterclockwise rotates in the $XY$ center slice. As $\theta_{\rm o}\to 0$, this central axis will align with the Y-axis.
In summary, the cohomogeneity–one Myers–Perry black hole ($a=b$) exhibits high degrees of symmetry in its hypershadow. The shape remains invariant under changes in the observer inclination angle $\theta_{\rm o}$, resulting only in rigid rotations. Additionally, the hypershadow demonstrates mirror symmetry with respect to the $XY$ plane and continuous rotational symmetry around the central axis, consistent with the enhanced Killing symmetries of the metric.
@@ 3.2.2 Rotation in a single plane $a=0$
In addition to the cohomogeneity–one case, we further investigate the single–rotation Myers–Perry black hole by setting $a=0$. This configuration is physically simpler, yet it lacks the additional Killing vectors that exist when $a=b$. As a result, it is difficult to derive an analytical shadow boundary. To overcome this limitation, we apply our novel numerical approach to compute the 3D hypershadow in this case. This method not only fills the analytical gap but also allows a detailed study of the shape and symmetry of the hypershadow under single plane rotation.
The line element for the single–rotation Myers–Perry black hole is given explicitly as follows[Emparan:2008eg]:
@@
$\Phi$ $=\mathbf{p}\,\sqrt{g_{\phi\phi}}\cos\alpha\sin\beta$ (3.20a)
$E$ $=\mathbf{p}\,(\frac{1+C^{t}\sqrt{g_{\psi\psi}}\cos\alpha\cos\beta\sin\zeta}{A^{t}})$ (3.20b)
Using the Hamilton equations[Diemer:2014lba, Kagramanova:2012hw, Bugden:2018uya] together with the initial conditions, we employ the backward ray–tracing method to obtain the 3D hypershadow.
@@
Figure 8: Hypershadows of the Myers–Perry black hole ($a=0,b=0.95$) rendered with different $N_{pix}$ and viewpoints. And the observer position (the initial condition) is located at $\{65,\pi/2-1/20\}$.
To further investigate the geometric structure of the hypershadow, we fix the observation angles $\{\alpha,\beta,\zeta\}$ to zero and compute cross–sectional slices through its center along each coordinate plane, in order to examine and discuss their symmetries.
@@
[…]
From Fig 9 we observe that the hypershadow exhibits mirror symmetry with respect to both the $XY$ plane (i.e., under $Z\mapsto-Z$) and the $YZ$ plane (i.e., under $X\mapsto-X$). These two mirror symmetry presist under different initial conditions.
Building on the discussion of continuous symmetry in the Myers–Perry black hole with equal angular momenta, we proceed to plot the difference maps under vertical reflection in the $XY$-plane for $\theta_{\rm o}=\pi/2-1/20$ and $\theta_{\rm o}=\pi/2-1/100$, in order to examine whether the single-rotation Myers–Perry black hole also exhibits continuous symmetry.
@@
[…]
From Fig. 10, we observe that as $\theta_{\rm o}\to\pi/2$, the difference between the original slice and its vertical reflection becomes negligible. This behavior closely resembles the $a=b$ case, indicating that as $\theta_{\rm o}\to\pi/2$, the hypershadow approaches a configuration exhibiting continuous rotational symmetry about the XX–axis.
@@
As $\theta_{\rm o}$ decreases toward zero, the central slice in the $XY$ plane undergoes a noticeable deformation. From Fig. 11, we observe that the hypershadow of the Myers–Perry black hole gradually approaches a nearly circular shape, closely resembling that of the Tangherlini black hole. However, its center simultaneously shifts away from the image center. These findings are further supported by the full three–dimensional hypershadow renderings presented below.
@@
Figure 12: Hypershadows of the Myers–Perry black hole ($a=0,b=0.95$) rendered with different $N_{\rm pix}$ and observer position. The resolutions is set to $N_{\rm pix}=64$ in the first row, $N_{\rm pix}=24$ in the second row.
In comparison, the hypershadow of the single–rotation Myers–Perry black hole ($a=0)$ exhibits a distinct behavior. While it retains mirror symmetry with respect to both the $XY$ and $YZ$ planes, its shape gradually transitions toward a more spherical structure as $\theta_{\rm o}$ decreases. Unlike the cohomogeneity–one case, where the hypershadow undergoes rigid rotation under changes in the observer inclination angle, the single–rotation configuration results in nontrivial shape deformation. These results highlight the richer geometric response in less symmetric spacetimes and demonstrate the flexibility of our numerical framework in capturing such variations.
@@ 4 Effects of Observer Position on the Black Hole Shadow
In this section we study how the observer position shapes the black hole shadow (hypershadow). It is well known that increasing the observer radial coordinate $r_{o}$ (denoted $x_{\rm o}=r_{o}^{2}$ in five dimensions) decreases the apparent size of the shadow. The influence of the observer inclination angle $\theta_{\rm obs}$ has been analyzed extensively in Ref. [Vazquez:2003zm, Papnoi:2014aaa, Hu:2020usx]. They work with angular (celestial) coordinates or impact-parameter mappings in more direct, conventional ways using a projection different from that adopted in our work. By contrast, in the equirectangular projection used in this paper, it has received little attention, with only a few studies[Novo:2024wyn] addressing it. Here we present a detailed analysis of this effect within the equirectangular projection.
To set the stage, we briefly review the $\theta_{\rm o}$–dependence in the four–dimensional Schwarzschild and Kerr cases; see Fig. 13. When $\theta_{\rm o}\to 0$, the shadows become circular. In the Kerr case, frame dragging near the boundary distorts the photon trajectories, while the Schwarzschild case retains symmetric deflection without dragging. These color–coded panels conveniently illustrate how light rays are gravitationally deflected and either captured by the black hole or escape to infinity, thereby outlining the shadow boundary.
@@
In higher dimensions such a visualization of individual light–ray trajectories is no longer feasible. Instead, the impact of $\theta_{\rm o}$ can be inferred from two image–space diagnostics of the hypershadow: (i) deformation (changes in shape) and (ii) displacement (a global offset of the image). The corresponding five–dimensional results are shown in Figs. 7 and 1112. In what follows we discuss the geometric origin of these features and then quantify the variations with observer inclination.
The physical origin can be traced to the two orthogonal rotation planes of the five-dimensional Myers–Perry geometry,
@@
where the aa–spin acts in the $\phi$–plane (weighted by $\sin^{2}\theta$) and the bb–spin in the $\psi$–plane (weighted by $\cos^{2}\theta$). Here $\theta$ is the $S^{3}$ angle controlling the local coupling of each spin, whereas $\theta_{\rm o}$ selects the mix sampled along the line of sight. In what follows we focus on two representative spin configurations that already exhibit sharply different responses:
(i) Cohomogeneity–one ($a=b$). The two planes contribute symmetrically; frame–dragging effects combine isotropically in the image space, so varying $\theta_{\rm o}$ produces a rigid rotation without displacement or distortion.
(ii) Single spin ($a=0$). In the singly rotating case ($a=0$) the two effects separate cleanly at the endpoints of $\theta_{\rm o}$. When the observer is aligned with the rotation axis ($\theta_{\rm o}=0$), frame dragging acts almost uniformly in the local mapping of impact parameters, producing a global offset while the hypershadow boundary remains nearly circular. By contrast, when the observer lies in the rotational equator ($\theta_{\rm o}=\pi/2$), the global offset vanishes and the boundary exhibits the strongest differential distortion. For intermediate angles $0<\theta_{\rm o}<\pi/2$, both contributions are present: the centroid (visual-center) shift gradually decreases with $\theta_{\rm o}$, whereas the boundary distortion increases, yielding a smooth trade-off between translation and deformation. Taken together, the $a=b$ case highlights the robustness of rotational symmetry, whereas the $a=0$ case shows that $\theta_{\rm o}$ is a sensitive probe of rotational anisotropy in higher dimensions.
@@
Cohomogeneity–one case ($a=b$) vs. Schwarzschild–Tangherlini.
We first compare the voxel counts of the cohomogeneity-one Myers–Perry hypershadow with those of the Schwarzschild–-Tangherlini (ST) black hole at the same $x_{\rm o}$ and resolution. The results show that all voxels of the MP hypershadow are included within the ST reference, with no additional points protruding beyond it. Accordingly, the total number of captured voxels in the MP case is smaller than in the ST case, indicating that rotation reduces the overall size of the hypershadow without producing any extra boundary features.
Single–spin case ($a=0$) vs. Schwarzschild–Tangherlini.
For the singly rotating Myers–Perry black hole, the hypershadow exhibits mirror symmetry with respect to both the $XY$ plane ($Z\mapsto-Z$) and the $YZ$ plane ($X\mapsto-X$), see Fig. 9. Due to this symmetry, the visual center can be determined simply as the midpoint between the top and bottom voxel along the YY–axis. We define the displacement $\eta$ as the distance of this midpoint from the origin. After shifting the ST reference by $\eta$ in the YY–direction, it still fully covers the MP hypershadow.
Taken together, the results indicate that black hole spin primarily shrinks the hypershadow and produces a global offset, without generating new boundary features, consistent with earlier observations [Papnoi:2014aaa]. To probe subtle deviations from spherical symmetry in the hypershadow, we introduce a distortion parameter $\delta_{s}$. This definition, appropriate for the 3D setting, differs from the 2D shadow distortion employed in Ref. [Papnoi:2014aaa].
We evaluate the total number of captured voxels at fixed radial coordinate $x_{\rm o}$ and resolution $N_{\rm pix}^{3}$. Let $N_{\rm cap}^{\rm ST}$ and $N_{\rm cap}^{\rm MP}$ denote the captured–voxel counts for the Schwarzschild–Tangherlini (ST) and Myers–Perry (MP) cases, respectively. We measure the deficit relative to the ST baseline by
@@
Table 1: Captured voxel statistics for the single–rotation Myers–Perry black hole.
@@
At the working resolution $64^{3}$, the Schwarzschild–Tangherlini (ST) baseline yields $N_{\rm cap}^{\rm ST}=35112$ captured voxels. For the singly rotating Myers–Perry case ($a=0$), Table. 1 summarizes, for each observer inclination and spin $b$, the captured-voxels counts $N_{\rm cap}^{\rm MP}$, the deficit $N_{\rm loss}$, and the distortion parameter $\delta_{s}$. Two clear trends emerge: (i) At fixed $\theta_{\rm o}$, $\delta_{s}$ decreases monotonically as $b$ is reduced ; (ii) At fixed $b$, $\delta_{s}$ grows monotonically as the observer inclination angle increase. Moreover, increasing $b$ enhances the angular contrast: for larger $b$ the distortion parameter is well described by a $\sin^{2}\!\theta_{\rm o}$ dependence, whereas for smaller $b$ the proportionality is less evident.
Table 2: Myers–Perry Captured voxel statistics, $\theta_{\rm o}=\pi/2-1/20$: $a=0$ vs. $a=b$.
@@
Table. 2 reports, for the near–equatorial view ($\theta_{\rm o}\approx\pi/2$), the captured-voxel count $N_{\rm cap}^{\rm MP}$, the deficit $N_{\rm loss}$, and the distortion parameter $\delta_{s}$ for both the cohomogeneity–one case ($a=b$) and the single–rotation case ($a=0$). From these data we infer that, at fixed spin $b$, the distortion in the cohomogeneity–one case is always greater than or equal to that in the single–rotation case.
Having established the size trend, we now turn to the displacement. Specifically, we evaluate the center shift (displacement) $\eta$ on the central $YZ$ slice ($X=0$) for the single–rotation configuration ($a=0$), using an in–slice resolution of $512^{2}$.
Table 3: Displacement of the hypershadow in the single–rotation case.
@@
From Table. 3, for the singly rotating Myers–Perry black hole ($a=0$) we find that the hypershadow displacement $\eta$ increases as the observer inclination decreases at fixed rotation parameter $b$, and, at fixed $\theta_{\rm o}$, grows monotonically with $b$.
Relative to the Schwarzschild–Tangherlini baseline, spin reduces the hypershadow size: the distortion parameter $\delta_{s}$ increases with $b$ in both configurations, and in the single rotation case ($a=0$) also grows with the observer inclination. Displacement is specific to $a=0$: the center shift $\eta$ increases with $b$ and strengthens as $\theta_{\rm o}$ decreases, whereas for $a=b$ the hypershadow undergoes a rigid rotation with no displacement.

### v3

@@ 3.2 Myers-Perry black hole
In 1986, Myers and Perry derived a stationary, axisymmetric, and asymptotically flat vacuum solution as a generalization of the Kerr black hole to higher dimensions[16]. The 5D Myers-Perry solution is given by the line element:
@@
The black hole mass is controlled by the same parameter $\mu$ introduced in Eq (3.2), and its two angular momenta are
@@
while $a$ and $b$ are the spin parameters associated with two independent rotation planes. These parameters are bounded by the condition $\mu\geq|a|+|b|$, which ensures the existence of an event horizon. The angular coordinates $\phi$ and $\psi$ range over the interval $0,2\pi]$, while $\theta$ takes values in $0,\pi/2]$.
@@
The metric (3.5) is invariant under the following transformation:
@@
$\cos\bar{\theta}\,\partial_{\bar{\theta}}-\cot\bar{\theta}\sin\bar{\phi}\,\partial_{\bar{\phi}}+\frac{\sin\bar{\phi}}{\sin\bar{\theta}}\,\partial_{\bar{\psi}}$ (3.10)
@@
$\sin\bar{\phi}\,\partial_{\bar{\theta}}-\cot\bar{\theta}\cos\bar{\phi}\,\partial_{\bar{\phi}}+\frac{\cos\bar{\phi}}{\sin\bar{\theta}}\,\partial_{\bar{\psi}},$ (3.11)
@@
3.2.1 Cohomogeneity-one solution: $a=b$
The shadow of the five-dimensional Myers-Perry black hole was first investigated through two-dimensional projection images by [19]. Recently, the work [18] provided an analytical description of the hypershadow boundary in the cohomogeneity-one case, with $a=b$. We compute the hypershadow numerically, investigate its physical features, and verify the consistency with the analytical results.
In order to study the cohomogeneity-one Myers-Perry spacetime ($a=b$), the line element can be written as [17]
@@ 3.2.1 Cohomogeneity-one solution: $a=b$
By evolving the geodesic equations[12, 8] derived from the Hamiltonian formalism under specified initial conditions, we reconstruct the 3D hypershadow using the backward ray-tracing method (see Fig. 4 ).
@@
Figure 4: Hypershadows of the Myers-Perry black hole ($a=b=0.45$) rendered with different $N_{\rm pix}$ and viewpoints. And the observer position is located at $\{65,\pi/2-1/20\}$.
To probe the geometric structure of the hypershadow, we fix the observation angles $\{\alpha,\beta,\zeta\}$ to zero and compute cross-sectional slices through its center along each coordinate plane, in order to examine and discuss their symmetries.
@@
Figure 5: Top row: central slices of the hypershadow in the $XY$, XZXZ, and $YZ$ planes. Middle row: H-$\Delta$ maps. Bottom row: V-$\Delta$ maps. Parameters: $a=b=0.45$, observer at $\{65,\pi/2-1/20\}$, and resolution $N_{\rm pix}=512$.
@@
In this paper [18], the authors proposed a simpler parametrization of the Myers-Perry (MP) hypershadow based on analytical construction. They concluded that the shape of the hypershadow is independent of the observer inclination angle $\theta_{\rm o}$. Varying $\theta_{\rm o}$ results only in a rigid rotation of the hypershadow in the image space, while the overall structure remains unchanged. Moreover, the geometry exhibits continuous rotational symmetry about the central axis. When $\theta_{\rm o}=\pi/2$, this central axis aligns with the $X$ axis.
@@
Figure 6: XYXY (H-$\Delta$) with different $\theta_{\rm o}$.
From Fig. 6, we find that as $\theta_{\rm o}\to\pi/2$ the difference between the original slice and its vertical reflection diminishes significantly. This suggests that the $XY$ central slice becomes symmetric under vertical reflection when $\theta_{\rm o}=\pi/2$. Consequently, this implies the existence of continuous rotational symmetry about the central axis. After confirming the approximate vertical symmetry near $\theta_{\rm o}\approx\pi/2$, we proceed to analyze the central slices in the $XY$ plane for various observer inclination angles, in order to further illustrate the rotational behavior of the hypershadow.
@@
Figure 7: XYXY central slices with different $\theta_{\rm o}$.
From Fig. 7, we could get that when we decrease the value of $\theta_{\rm o}$ the central axis will counterclockwise rotates in the $XY$ center slice. As $\theta_{\rm o}\to 0$, this central axis will align with the $Y$ axis.
In summary, the cohomogeneity-one Myers-Perry black hole ($a=b$) exhibits high degrees of symmetry in its hypershadow. The shape remains invariant under changes in the observer inclination angle $\theta_{\rm o}$, resulting only in rigid rotations. Additionally, the hypershadow demonstrates mirror symmetry with respect to the $XY$ plane and continuous rotational symmetry around the central axis, consistent with the enhanced Killing symmetries of the metric.
@@ 3.2.2 Rotation in a single plane $a=0$
In addition to the cohomogeneity-one case, we further investigate the single-rotation Myers-Perry black hole by setting $a=0$. This configuration is physically simpler, yet it lacks the additional Killing vectors that exist when $a=b$. As a result, it is difficult to derive an analytical shadow boundary. To overcome this limitation, we apply our novel numerical approach to compute the 3D hypershadow in this case. This method not only fills the analytical gap, but also allows a detailed study of the shape and symmetry of the hypershadow under single plane rotation.
The line element for the single-rotation Myers-Perry black hole is given explicitly as follows[9]:
@@
$\Phi$ $=\mathbf{p}\,\sqrt{g_{\phi\phi}}\cos\alpha\sin\beta,$ (3.20a)
$E$ $=\mathbf{p}\,(\frac{1+C^{t}\sqrt{g_{\psi\psi}}\cos\alpha\cos\beta\sin\zeta}{A^{t}}).$ (3.20b)
Using the Hamilton equations[8, 14, 4] together with the initial conditions, we employ the backward ray-tracing method to obtain the 3D hypershadow (see Fig. 8).
@@
Figure 8: Hypershadows of the Myers-Perry black hole ($a=0,b=0.95$) rendered with different $N_{\rm pix}$ and viewpoints. And the observer position (the initial condition) is located at $\{65,\pi/2-1/20\}$.
To further investigate the geometric structure of the hypershadow, we fix the observation angles $\{\alpha,\beta,\zeta\}$ to zero and compute cross-sectional slices through its center along each coordinate plane, in order to examine and discuss their symmetries.
@@
Figure 9: Top row: central slices of the hypershadow in the $XY$, XZXZ, and $YZ$ planes. Middle row: H-$\Delta$ maps. Bottom row: V-$\Delta$ maps. Parameters: singly rotating case ($a=0$, $b=0.95$), observer at $\{65,\pi/2-1/20\}$, and resolution $N_{\rm pix}=512$.
From Fig. 9, we observe that the hypershadow exhibits mirror symmetry with respect to both the $XY$ plane (i.e., under $Z\mapsto-Z$) and the $YZ$ plane (i.e., under $X\mapsto-X$). These two mirror symmetries presist under different initial conditions.
Building on the discussion of continuous symmetry in the Myers-Perry black hole with equal angular momenta, we proceed to plot the difference maps under vertical reflection in the $XY$ plane for $\theta_{\rm o}=\pi/2-1/20$ and $\theta_{\rm o}=\pi/2-1/100$, in order to examine whether the single-rotation Myers-Perry black hole also exhibits continuous symmetry.
@@
Figure 10: XYXY (H-$\Delta$) with different $\theta_{\rm o}$.
From Fig. 10, we observe that as $\theta_{\rm o}\to\pi/2$ the difference between the original slice and its vertical reflection becomes negligible. This behavior closely resembles the $a=b$ case, indicating that as $\theta_{\rm o}\to\pi/2$ the hypershadow approaches a configuration exhibiting continuous rotational symmetry about the $X$ axis.
@@
As $\theta_{\rm o}$ decreases toward zero, the central slice in the $XY$ plane undergoes a noticeable deformation. From Fig. 11, we observe that the hypershadow of the Myers-Perry black hole gradually approaches a nearly circular shape, closely resembling that of the Tangherlini black hole. However, its center simultaneously shifts away from the image center. These findings are further supported by the full three-dimensional hypershadow renderings presented in Fig. 12.
@@
Figure 12: Hypershadows of the Myers-Perry black hole ($a=0,b=0.95$) rendered with different $N_{\rm pix}$ and observer position. The resolutions is set to $N_{\rm pix}=64$ in the first row and $N_{\rm pix}=24$ in the second row.
In comparison, the hypershadow of the single-rotation Myers-Perry black hole ($a=0)$ exhibits a distinct behavior. While it retains mirror symmetry with respect to both the $XY$ and $YZ$ planes, its shape gradually transitions toward a more spherical structure as $\theta_{\rm o}$ decreases. Unlike the cohomogeneity-one case, where the hypershadow undergoes rigid rotation under changes in the observer inclination angle, the single-rotation configuration results in nontrivial shape deformation. These results highlight the richer geometric response in less symmetric spacetimes and demonstrate the flexibility of our numerical framework in capturing such variations.
@@ 4 EFFECTS OF OBSERVER POSITION ON THE BLACK HOLE SHADOW
In this section, we study how the observer position shapes the black hole shadow (hypershadow). It is well known that increasing the observer radial coordinate $r_{o}$ (denoted $x_{\rm o}=r_{o}^{2}$ in five dimensions) decreases the apparent size of the shadow. The influence of the observer inclination angle $\theta_{\rm obs}$ has been analyzed extensively in Refs. [25, 19, 13]. They work with angular (celestial) coordinates or impact-parameter mappings in more direct, conventional ways using a projection different from that adopted in our work. By contrast, in the equirectangular projection used in this paper, it has received little attention, with only a few studies[18] addressing it. Here we present a detailed analysis of this effect within the equirectangular projection.
To set the stage, we briefly review the $\theta_{\rm o}$ dependence in the four-dimensional Schwarzschild and Kerr cases; see Fig. 13. When $\theta_{\rm o}\to 0$, the shadows become circular. In the Kerr case, frame dragging near the boundary distorts the photon trajectories, while the Schwarzschild case retains symmetric deflection without dragging. These color-coded panels conveniently illustrate how light rays are gravitationally deflected and either are captured by the black hole or escape to infinity, thereby outlining the shadow boundary.
@@
In higher dimensions such a visualization of individual light-ray trajectories is no longer feasible. Instead, the impact of $\theta_{\rm o}$ can be inferred from two image-space diagnostics of the hypershadow: (i) deformation (changes in shape) and (ii) displacement (a global offset of the image). The corresponding five-dimensional results are shown in Figs. 7, 11, and12. In what follows we discuss the geometric origin of these features and then quantify the variations with observer inclination.
The physical origin can be traced to the two orthogonal rotation planes of the five-dimensional Myers-Perry geometry:
@@
where the $a$ spin acts in the $\phi$ plane (weighted by $\sin^{2}\theta$) and the $b$ spin in the $\psi$ plane (weighted by $\cos^{2}\theta$). Here, $\theta$ is the $S^{3}$ angle controlling the local coupling of each spin, whereas $\theta_{\rm o}$ selects the mix sampled along the line of sight. In what follows, we focus on two representative spin configurations that already exhibit sharply different responses.
(i) Cohomogeneity-one ($a=b$). The two planes contribute symmetrically; frame-dragging effects combine isotropically in the image space, so varying $\theta_{\rm o}$ produces a rigid rotation without displacement or distortion.
(ii) Single spin ($a=0$). In the singly rotating case ($a=0$), the two effects separate cleanly at the end points of $\theta_{\rm o}$. When the observer is aligned with the rotation axis ($\theta_{\rm o}=0$), frame dragging acts almost uniformly in the local mapping of impact parameters, producing a global offset while the hypershadow boundary remains nearly circular. By contrast, when the observer lies in the rotational equator ($\theta_{\rm o}=\pi/2$), the global offset vanishes and the boundary exhibits the strongest differential distortion. For intermediate angles $0<\theta_{\rm o}<\pi/2$, both contributions are present: The centroid (visual-center) shift gradually decreases with $\theta_{\rm o}$, whereas the boundary distortion increases, yielding a smooth trade-off between translation and deformation. Taken together, the $a=b$ case highlights the robustness of rotational symmetry, whereas the $a=0$ case shows that $\theta_{\rm o}$ is a sensitive probe of rotational anisotropy in higher dimensions.
@@
Cohomogeneity-one case ($a=b$) vs. Schwarzschild-Tangherlini.
We first compare the voxel counts of the cohomogeneity-one Myers-Perry hypershadow with those of the Schwarzschild-Tangherlini (ST) black hole at the same $x_{\rm o}$ and resolution. The results show that all voxels of the MP hypershadow are included within the ST reference, with no additional points protruding beyond it. Accordingly, the total number of captured voxels in the MP case is smaller than in the ST case, indicating that rotation reduces the overall size of the hypershadow without producing any extra boundary features.
Single-spin case ($a=0$) vs. Schwarzschild-Tangherlini.
For the singly rotating Myers-Perry black hole, the hypershadow exhibits mirror symmetry with respect to both the $XY$ plane ($Z\mapsto-Z$) and the $YZ$ plane ($X\mapsto-X$); see Fig. 9. Because of this symmetry, the visual center can be determined simply as the midpoint between the top and bottom voxel along the $Y$ axis. We define the displacement $\eta$ as the distance of this midpoint from the origin. After shifting the ST reference by $\eta$ in the $Y$ direction, it still fully covers the MP hypershadow.
Taken together, the results indicate that black hole spin primarily shrinks the hypershadow and produces a global offset, without generating new boundary features, consistent with earlier observations [19]. To probe subtle deviations from spherical symmetry in the hypershadow, we introduce a distortion parameter $\delta_{s}$. This definition, appropriate for the 3D setting, differs from the 2D shadow distortion employed in Ref. [19].
We evaluate the total number of captured voxels at fixed radial coordinate $x_{\rm o}$ and resolution $N_{\rm pix}^{3}$. Let $N_{\rm cap}^{\rm ST}$ and $N_{\rm cap}^{\rm MP}$ denote the captured-voxel counts for the ST and MP cases, respectively. We measure the deficit relative to the ST baseline by
@@
Table 1: Captured voxel statistics for the single-rotation Myers-Perry black hole.
@@
At the working resolution $64^{3}$, the ST baseline yields $N_{\rm cap}^{\rm ST}=35112$ captured voxels. For the singly rotating Myers-Perry case ($a=0$), Table. 1 summarizes, for each observer inclination and spin $b$, the captured-voxel counts $N_{\rm cap}^{\rm MP}$, the deficit $N_{\rm loss}$, and the distortion parameter $\delta_{s}$. Two clear trends emerge: (i) At fixed $\theta_{\rm o}$, $\delta_{s}$ decreases monotonically as $b$ is reduced ; (ii) at fixed $b$, $\delta_{s}$ grows monotonically as the observer inclination angle increases. Moreover, increasing $b$ enhances the angular contrast: For larger $b$ the distortion parameter is well described by a $\sin^{2}\,\theta_{\rm o}$ dependence, whereas for smaller $b$ the proportionality is less evident.
Table 2: Myers-Perry captured-voxel statistics, $\theta_{\rm o}=\pi/2-1/20$: $a=0$ vs. $a=b$.
@@
Table. 2 reports, for the near-equatorial view ($\theta_{\rm o}\approx\pi/2$), the captured-voxel count $N_{\rm cap}^{\rm MP}$, the deficit $N_{\rm loss}$, and the distortion parameter $\delta_{s}$ for both the cohomogeneity–one case ($a=b$) and the single-rotation case ($a=0$). From these data we infer that, at fixed spin $b$, the distortion in the cohomogeneity-one case is always greater than or equal to that in the single-rotation case.
Having established the size trend, we now turn to the displacement. Specifically, we evaluate the center shift (displacement) $\eta$ on the central $YZ$ slice ($X=0$) for the single-rotation configuration ($a=0$), using an in-slice resolution of $512^{2}$.
Table 3: Displacement of the hypershadow in the single-rotation case.
@@
From Table 3, for the singly rotating Myers-Perry black hole ($a=0$), we find that the hypershadow displacement $\eta$ increases as the observer inclination decreases at fixed rotation parameter $b$ and, at fixed $\theta_{\rm o}$, grows monotonically with $b$.
Relative to the Schwarzschild-Tangherlini baseline, spin reduces the hypershadow size: The distortion parameter $\delta_{s}$ increases with $b$ in both configurations and in the single-rotation case ($a=0$) also grows with the observer inclination. Displacement is specific to $a=0$: The center shift $\eta$ increases with $b$ and strengthens as $\theta_{\rm o}$ decreases, whereas for $a=b$ the hypershadow undergoes a rigid rotation with no displacement.